Hello,
The answer has been provided, I am having dificulties showing how to arrive at the answers given. Can you please show how the problems are broken down for the desinated answers. Thank you

pg. 451
54) x+3y=2
-x+y=1
answer: (-1/4,3/4)


pg. 460
34)
3/7x + 5/9y =27
1/9x +2/7y =7
answer: (63,0)



pg. 461
50)
2y-x=3
x=3y-5
answer: (1,2)

pg. 510 Ch.8
18) -3x>0 and 3x-4 < 11

pg. 510 Ch.8
32) x > 3 or x < -3 ___-5____-4__-3__-2
__-1___0___1___2___3____4____5

)----------------------------------------(

pg 510 Ch. 8
50) -1 (< greater than or equal to) 3-2x<11
___-5____-4__-3__-2 __-1___0___1___2___3____4____5

(----------------------------------------]
30) 3 -1/2 absolute value l 1/2x - 4 l=2


62) 1/3 absolute value 2x-4 l < 1
___-5____-4__-3__-2 __-1___0___1___2___3____4____5

(--------------------)
pg. 521 Ch. 8
word problem
86) In July 1985 steve Cram of Great Britain ser a world record of 3
minutes 29.67 seconds for the 1500-meter race and a world record of 3
minute 46.31 seconds for the 1-miles race ( The Doubleday Almanac).
Suppose you know only that these teo events occurred 11 days apart and
that the 1500-meter record was set on July 16. Use an absolute value
equation to find the possible dates for the 1-miles record run.

pg. 529 Ch. 8
18) X > equal to or greater than -2 and x - 3y < 6

Answer by stanbon(75887)

pg. 451
54)
x+3y=2
-x+y=1
----------
Add the equations to get:
4y=3
y = 3/4
---------
Substitute to solve for x"
-x+(3/4) = 1
x = -1/4
answer: (-1/4,3/4)
=======================

pg. 460
34)
3/7x + 5/9y =27
1/9x +2/7y =7
--------------
Multiply thru 1st equation by 63; Multiply thru 2nd eqution by 63:
27x + 35y = 27*63
7x+18y = 7*63
-----------------
Multiply thru the 2nd equation by (27/7) to get:
27x + 35y = 27*63
27x+(27*18/7)y = 27*63
-------------
Subtract the 2nd equation from the 1st and solve for y to get:
y = 0
---
Substitute to solve for "x":
27x + 0 = 27*63
x = 63
answer: (63,0)
-------------------------


pg. 461
50)
2y-x=3
x=3y-5
-------------
Substitute to solve for "y":
2y-(3y-5)= 3
-y=-2
y = 2
--------
Substitute to solve for "x":
x= 3*2-5 = 1

answer: (1,2)
-------------------------
pg. 510 Ch.8
18) -3x>0 and 3x-4 < 11
x < 0 and 3x <15
x<0 and x < 5
-----------
Only x<0 satisfies both of these conditions.
--------------------
pg. 510 Ch.8
32) x > 3 or x < -3 ___-5____-4__-3__-2
__-1___0___1___2___3____4____5
)----------------------------------------(
pg 510 Ch. 8
50) -1 (< greater than or equal to) 3-2x<11
___-5____-4__-3__-2 __-1___0___1___2___3____4____5
(----------------------------------------]
30) 3 -1/2 absolute value l 1/2x - 4 l=2

62) 1/3 absolute value 2x-4 l < 1
___-5____-4__-3__-2 __-1___0___1___2___3____4____5
(--------------------)
pg. 521 Ch. 8
word problem
86) In July 1985 steve Cram of Great Britain set a world record of 3
minutes 29.67 seconds for the 1500-meter race and a world record of 3
minute 46.31 seconds for the 1-miles race ( The Doubleday Almanac).
Suppose you know only that these two events occurred 11 days apart and
that the 1500-meter record was set on July 16. Use an absolute value
equation to find the possible dates for the 1-miles record run.
Let the date of the 1-mile record run be "x".
INEQUALITY:
5<=x<=27
|x-16|<=11
-----------------
pg. 529 Ch. 8
18) X > equal to or greater than -2 and x - 3y < 6
===================
Cheers,
Stan H.